### v1

# A look on equations describing pseudospherical surfaces

Abstract
We revisit the notion of equations describing pseudospherical surfaces starting from its roots influenced by the AKNS system until current research topics in the field.
Keywords Equations describing pseudospherical surfaces  $\cdot$ Camassa-Holm equation
Dedicatory: To Professor Keti Tenenblat on occasion of her 80th birthday.
1 Introduction
In [1] the authors considered the following linear system
$\partial_{x}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\begin{pmatrix}-i\zeta&q\\ \\ r&i\zeta\end{pmatrix}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix},$ (1.0.1)
$\partial_{t}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\begin{pmatrix}A&B\\ \\ C&D\end{pmatrix}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix},$ (1.0.2)
where, $q$ and $r$ are functions of $(x,t)$ whereas $A$, $B$, $C$ and $D$ depend on $x,\,t$ and $\zeta$, which may eventually depend on t. The field variables $x$ and $t$ can be regarded as space and time, respectively.
Equation (1.0.1) is a scattering problem while (1.0.2) is an evolution problem.
From the cross differentiation and the integrability condition
$\partial_{t}\partial_{x}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\partial_{x}\partial_{t}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}$
(1.0.1) and (1.0.2) give
$\begin{pmatrix}\partial_{x}A&\partial_{x}B\\ \\ \partial_{x}C&\partial_{x}D\end{pmatrix}-\begin{pmatrix}0&\partial_{t}q\\ \\ \partial_{t}r&0\end{pmatrix}=i\partial_{t}\zeta\begin{pmatrix}-1&0\\ \\ 0&1\end{pmatrix}+\begin{pmatrix}qC-rB&q(D-A)-2i\zeta B\\ \\ r(A-D)+2i\zeta C&rB-qC\end{pmatrix}.$
The $2\times 2$ matrix in (1.0.1) is traceless, as are all but the first in the above equation. If we impose that the second matrix in (1.0.2) be traceless, then all matrices we have so far will share this property. Without loss of generality, we may assume $D=-A$.
In addition, if we assume that the parameter $\zeta$ is isospectral, that is, it is independent of time, then we get the system of PDEs
$\left\{\begin{array}[]{lcl}\partial_{x}A&=&qC-rB,\\ \\ \partial_{x}B+2i\zeta B&=&\partial_{t}q-2Aq,\\ \\ \partial_{x}C-2i\zeta C&=&\partial_{t}r+2Ar.\end{array}\right.$ (1.0.3)
The reader may be wondering: What is the relevance of (1.0.3)? This is a very good question.
In [1] it was shown that a number of very important models, such as the KdV and the sine-Gordon equations, can be obtained from (1.0.3) depending on the choices of the involved functions. Moreover, equations arising from the compatibility condition (1.0.3) are said to be integrable (à la AKNS).
Let us give two examples illustrating how equations can be obtained from (1.0.3). The first one is a very classical case, while the second one is focused on a shallow water model unknown ty the time [1] was published.
Example 1.1.
Consider the choice
$A=\frac{i\cos{u}}{4},\,\,r=-q=\frac{u_{x}}{2},\,\,B=C=\frac{i\sin{u}}{4\zeta}.$
Then (1.0.3) is equivalent to the sine-Gordon (sG) equation
$u_{tx}=\sin{u}.$ (1.0.4)
It is worth mentioning that the sG equation was the first PSS equation to be discovered, even before the concept was brought to light. This comes from the fact that the sG plays vital a vital role in the context of the geometry of surfaces with constant and negative curvatures [36, Chapter 1].
Example 1.2.
Consider
$\begin{array}[]{l}A=\displaystyle{\frac{1}{2}\Big{(}u_{x}-\eta u-\frac{\beta}{\eta}\Big{)},\,\,\,B=\frac{1}{2}\Big{(}\beta u-1+\frac{\beta u}{\eta^{2}}+u+\frac{u}{\eta^{2}}\Big{)}}\\ \\ C=\displaystyle{\frac{1}{2}\Big{(}\frac{2}{\eta^{2}}-2um+\beta u+\frac{\beta u}{\eta^{2}}-u-2\frac{\beta u_{x}}{\eta}+2\frac{u_{x}}{\eta}-u-\frac{2\beta+u}{\eta^{2}}\Big{)}},\\ \\ r=\displaystyle{\frac{1}{2}\Big{(}1-2m-\beta+\frac{\beta+1}{\eta^{2}}\Big{)},\,\,q=\frac{1}{2}\Big{(}\frac{\beta-1}{\eta^{2}}-1-\beta\Big{)},\,\,m=u-u_{xx},\,\,\eta=-2i\zeta},\end{array}$
where the parameters satisfy the constraint
$\eta^{4}-\eta^{2}+\beta^{2}\eta^{2}=\beta^{2}+1-2\beta.$
Then (1.0.3) is equivalent to the Camassa-Holm (CH) equation
$u_{t}-u_{txx}+3uu_{x}=2u_{x}u_{xx}+uu_{xxx},$ (1.0.5)
or its equivalent form
$m_{t}+2um_{x}+u_{x}m=0.$ (1.0.6)
Let us now introduce a geometric view on (1.0.1)–(1.0.2). By $\{dx,dt\}$ we denote the usual dual basis of $\mathbb{R}^{2}$. Consider eight one-forms $\Omega_{ij},\,\Psi_{ij}$, $1\leq i,j\leq 2$ and define one-form matrices $\Omega=(\Omega_{ij})$, $\Psi=(\Psi_{ij})$. If $d$ denotes the usual exterior derivative, then we have
$d\Omega=(d\Omega_{ij})$
and we can define the product of two one-form matrices in a natural way by putting
$\Omega\wedge\Psi=\Big{(}\sum_{k=1}^{2}\Omega_{ik}\wedge\Psi_{kj}\Big{)}.$
All that said, let us consider the following one-form matrix
$\Omega=\begin{pmatrix}-i\zeta dx+Adt&qdx+Bdt\\ \\ rdx+Cdt&i\zeta dx-Adt\end{pmatrix}$ (1.0.7)
and consider
$dv=\Omega v,$ (1.0.8)
where $v=(v_{1}\quad v_{2})^{t}$ (the superscript $t$ means the usual matrix transposition). From (1.0.6) we obtain
$0=d(dv)=d(\Omega v)=(d\Omega-\Omega\wedge\Omega)v,$
meaning that
$d\Omega=\Omega\wedge\Omega.$ (1.0.9)
If we now define two other matrices $X$ and $T$, such that
$\Omega=Xdx+Tdt,$
then (1.0.7) gives the equation
$\partial_{t}X+\partial_{x}T+[X,T]=0.$ (1.0.10)
Equation (1.0.8) is better known as the zero curvature representation, and it reveals to us that (1.0.1)–(1.0.2) can be seen as a geometric object.
Equations (1.0.7) are a different, but equivalent, way to write (1.0.1)–(1.0.2). whereas (1.0.9) implies to (1.0.3).
The remainder of the paper is as follows: in Section 2 we revisit the main aspects of the works by Sasaki [38] and Chern and Tenenblat [8], connecting equations obtained from the AKNS method and the geometry of abstract 2-manifolds with constant and negative Gaussian curvature. In Section 3 we discuss some recent results from the author [16, 19, 20, 21, 37] that connect the theory developed by Chern and Tenenblat with problems concerning not only equations, but equations subject to some condition (initial data). In some parts, the presentation may closely follow the original references.
2 Equations describing pseudospherical surfaces
Chern and Tenenblat pioneered the concept of pseudospherical equations [8], establishing a connection between certain partial differential equations and smooth two-dimensional Riemannian manifolds. Their contribution gained particular significance following a notable insight by Sasaki, who had earlier demonstrated that solutions to specific integrable equations—those arising from the AKNS formalism [1]—are intimately related to the intrinsic geometry of surfaces with constant negative Gaussian curvature ${\cal K}=-1$ [38].
Definition 2.1.
A pseudospherical surface (PSS) is a two-dimensional Riemannian manifold whose Gaussian curvature is constant and negative.
These foundational results laid the groundwork for a new intersection between surface theory and the analysis of PDEs. Since then, this perspective has shaped a substantial body of literature, including [2, 5, 6, 7, 16, 19, 31, 32, 33, 34, 35, 37], among many others. A more comprehensive introduction to the topic may be found in [39, Chapter 1].
By definition, a pseudospherical surface (abbreviated as PSS, also used in the plural) is a two-dimensional Riemannian manifold whose Gaussian curvature remains constant and negative. According to Hilbert’s theorem [15, Section 5-11], such manifolds cannot be isometrically immersed in $\mathbb{R}^{3}$ as complete surfaces, due to their negative curvature—see also [28, page 439] and [27]. For this reason, they are frequently referred to as abstract surfaces in the literature.
Once these notions are introduced, let us precisely give meaning to them à la Chern and Tenenblat [8]. Let $(x,t)$ be independent variables. A differential equation for a real-valued function $u=u(x,t)$ of order $n$ is generically denoted by
${\cal E}(x,t,u,u_{(1)},\cdots,u_{(n)})=0.$ (2.0.1)
Definition 2.2.
A differential equation (2.0.1) is said to describe a pseudospherical surface, or it is said to be of pseudospherical type, if it is a necessary and sufficient condition for the existence of differentiable functions $f_{ij}$, $1\leq i,j\leq 3$, such that the forms
$\omega_{i}=f_{i1}dx+f_{i2}dt,\quad 1\leq i\leq 3,$ (2.0.2)
satisfy the structure equations of a pseudospherical surface
$d\omega_{1}=\omega_{3}\wedge\omega_{2},\quad d\omega_{2}=\omega_{1}\wedge\omega_{3},\quad d\omega_{3}=\omega_{1}\wedge\omega_{2}.$ (2.0.3)
Sasaki noticed in his work [38] that by defining one-forms $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$ suitably from (1.0.1) and (1.0.2), one could obtain the one-form matrix
$\Omega=\frac{1}{2}\begin{pmatrix}\omega_{2}&\omega_{1}-\omega_{3}\\ \omega_{1}-\omega_{3}&-\omega_{2}\end{pmatrix},$ (2.0.4)
$\Sigma:=d\Omega-\Omega\wedge\Omega.$ (2.0.5)
Then $\Sigma=0$ if and only if (2.0.3) holds. As a result, if (2.0.1) is the compatibility condition of (1.0.3), then its solutions satisfy (2.0.3), that is, the structure equations for a two-dimensional Riemannian manifold with first fundamental form, or metric, given by
$I=\omega_{1}^{2}+\omega_{2}^{2}.$ (2.0.6)
In practical terms, given a triad of one-forms $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$, and an equation (2.0.1), we can check if they describe a PSS surface in the following way: let us define a matrix of one-forms $\Omega$ by (2.0.4). If, when restricted to the manifold determined by the solutions of (2.0.1), the matrix $\Sigma$ vanishes, we say that (2.0.1) is a PSS equation, and the triad $\{\omega_{1},\,\omega_{2},\,\omega_{3}\}$ satisfies the structure equations (2.0.3) of a PSS with Gaussian curvature ${\cal K}=-1$. As a result, any open, simply connected set $U$ where the one-forms $\omega_{1}$ and $\omega_{2}$ are defined and $\omega_{1}\wedge\omega_{2}\neq 0$ everywhere, is endowed with a PSS structure. We can naively see this condition as a requirement for $\{\omega_{1},\omega_{2}\}$ to be a linearly independent set and thus be the basis of a plane.
Definition 2.3.
Assume that (2.0.1) is an equation describing a PSS and let $u$ be a solution of it. We say that $u$ is a generic solution if the one-forms (2.0.2) satisfy the condition $\omega_{1}\wedge\omega_{2}\neq 0$.
Example 2.1.
Let us consider the following one-forms
$\left\{\begin{array}[]{lcl}\omega_{1}&=&\displaystyle{\frac{1}{\eta}\sin{u}dt,}\\ \\ \omega_{2}&=&\displaystyle{\eta dx+\frac{1}{\eta}\cos{u}dt},\\ \\ \omega_{3}&=&u_{x}dx.\end{array}\right.$ (2.0.7)
Then we have
$\left\{\begin{array}[]{lcl}d\omega_{1}-\omega_{3}\wedge\omega_{2}&\equiv&\displaystyle{0}\\ \\ d\omega_{2}-\omega_{1}\wedge\omega_{3}&\equiv&\displaystyle{0},\\ \\ d\omega_{3}-\omega_{1}\wedge\omega_{2}&=&-(u_{tx}-\sin{u})dx\wedge dt.\end{array}\right.$
As a result, on the solutions of the sG equation (1.0.4), the forms (2.0.7) satisfy (2.0.3). Since
$\omega_{1}\wedge\omega_{2}=-\sin{u}\,dx\wedge dt,$
then $u$ will be a generic solution for (1.0.4) provided that $u\notin\{k\pi,\,\,\,k\in\mathbb{Z}\}$.
The generic solutions provide the first fundamental form
$I=\eta^{2}dx^{2}+2\cos{u}dxdt+\frac{1}{\eta^{2}}dt^{2}.$
Example 2.2.
Let $m=u-u_{xx}$, $\lambda\in\mathbb{R}\setminus\{0\}$ and consider the one-forms
$\begin{array}[]{lcl}\omega_{1}&=&\displaystyle{\Big{(}\frac{\lambda}{2}+\frac{1}{2\lambda}-m\Big{)}dx+\Big{(}um+\frac{\lambda}{2}u-\frac{u}{2\lambda}-\frac{1}{2}-\frac{\lambda^{2}}{2}\Big{)}dt},\\ \\ \omega_{2}&=&-u_{x}dt,\\ \\ \omega_{3}&=&\displaystyle{\Big{(}m+\frac{1}{2\lambda}-\frac{\lambda}{2}\Big{)}dx+\Big{(}\frac{\lambda^{2}}{2}-\frac{1}{2}-\frac{u}{2\lambda}-\frac{\lambda}{2}u-um\Big{)}dt}.\end{array}$ (2.0.8)
A straightforward calculation shows that
$\begin{array}[]{lcl}d\omega_{1}-\omega_{3}\wedge\omega_{2}&=&\Big{(}u_{t}-u_{txx}+3uu_{x}-2u_{x}u_{xx}-uu_{xxx}\Big{)}dx\wedge dt,\\ \\ d\omega_{2}-\omega_{1}\wedge\omega_{3}&=&0,\\ \\ d\omega_{3}-\omega_{1}\wedge\omega_{2}&=&-\Big{(}u_{t}-u_{txx}+3uu_{x}-2u_{x}u_{xx}-uu_{xxx}\Big{)}dx\wedge dt,\end{array}$ (2.0.9)
meaning that on the generic solutions of the CH equation (1.0.5), (2.0.8) define a two-dimensional Riemannian manifold with first fundamental form
$\begin{array}[]{lcl}I&=&\displaystyle{\Big{[}m^{2}-\Big{(}\lambda+\frac{1}{\lambda}\Big{)}m+\frac{1}{4}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}\Big{]}dx^{2}}+\\ \\ &&\displaystyle{\Big{[}-um^{2}+\frac{1}{\lambda}um+\frac{1}{4}\Big{(}\lambda^{2}-\frac{1}{\lambda^{2}}\Big{)}u+\lambda\Big{(}\frac{\lambda}{2}+\frac{1}{2\lambda}\Big{)}m-\lambda\Big{(}\frac{\lambda}{2}+\frac{1}{2\lambda}\Big{)}^{2}\Big{]}dxdt}+\\ \\ &&\displaystyle{\Big{[}um^{2}+u^{2}m\Big{(}\lambda-\frac{1}{\lambda}\Big{)}+u_{x}^{2}-\lambda\Big{(}\lambda+\frac{1}{\lambda}\Big{)}um+\frac{1}{4}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}u^{2}-\frac{\lambda}{2}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}}\\ \\ &&\displaystyle{+\frac{\lambda^{2}}{4}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}\Big{]}dt^{2}}.\end{array}$
Let us investigate the generic solutions of the CH equation. It is straightforward to see that
$\omega_{1}\wedge\omega_{2}=-\Big{(}\frac{\lambda}{2}+\frac{1}{2\lambda}-m\Big{)}u_{x}dx\wedge dt.$
Provided that $u$ is a solution of the CH equation, then $\omega_{1}\wedge\omega_{2}=0$ somewhere if and only if
$m=\frac{\lambda}{2}+\frac{1}{2\lambda}$ (2.0.10)
or
$u_{x}=0.$ (2.0.11)
We close the present section with some remarks.
Remark 2.1.
It is possible that a PSS equation possesses more than one representation of the forms satisfying the structure equations (2.0.3). If we substitute the choices made in Example 1.2 into (1.0.7) will lead to different one-forms (see [32, Theorem 1]) than those in (2.0.8) for the CH equation, although (1.0.10) would be equivalent to the CH equation (1.0.5). However, any of the triads given either in [32, Theorem 1] or Example 1.2 are enough to ensure that the CH equation describes PSS.
Remark 2.2.
The theory of PSS equations, even though purely geometric, was born in the context of integrable systems. That is why the earlier papers in the area were concerned with equations that are integrable. The intense research carried out by K. Tenenblat and her co-workers shows that the land of PSS equations is larger than that of integrable equations in the AKNS sense.
Remark 2.3.
The theory of PSS equations was conceived only for equations, or even systems of equations, in mind. Additional information, such as boundary or initial conditions, was not considered for a long time.
Remark 2.4.
The theory of PSS equations has traditionally been formulated under the assumption that the solutions defining the first fundamental form possess $C^{\infty}$ regularity. Although this smoothness requirement is not always explicitly stated, it can be traced back to the foundational work of Chern and Tenenblat [8, page 55, second paragraph]. In contrast, the discussion in the preceding paragraph highlights a tension between this framework and the class of solutions considered in [10, 11, 13], where the smoothness assumption is incompatible with the finite regularity typically associated with Cauchy problems. Consequently, surfaces potentially generated by solutions exhibiting wave breaking remained unexplored within this geometric context until [20].
3 Finite regularity of the metric
In this section we address problems concerning PSS equations and some developments made by the author in [16, 19, 20, 21, 22, 37]. The CH equation was the main motivation for such improvements and thus, parts of the text are now more focused on its literature. For a review on some aspects of the CH equation, see [18].
A smooth solution of a PSS equation yields smooth one-forms $\omega_{1},\omega_{2},\omega_{3}$, and consequently, the associated first fundamental form inherits the same degree of regularity. In contrast, the solutions considered by Constantin [13] are not necessarily $C^{\infty}$, which marks a significant distinction from works such as [6, 5, 31, 32, 33, 34, 35, 38], where smoothness is a standing assumption.
In the literature on PDEs and PSS structures, the uniqueness of solutions is not often addressed. Consequently, whether a specific first fundamental form may correspond to one or multiple solutions of PSS equations remains, overall, an open topic.
This distinction motivates a deeper investigation into the relationship between regular curves and the graphs of solutions of the CH equation. More precisely, one may ask under which conditions a regular curve in space can uniquely determine such a graph, together with a compatible set of one-forms. In certain cases, uniqueness can indeed be established. This situation essentially corresponds to solving the CH equation with prescribed initial data $u_{0}$, which naturally determines a curve of the form $(x,0,u_{0}(x))$. However, the requirement of uniqueness imposes constraints on the class of curves that can be considered.
As previously noted, solutions to Cauchy problems for the CH equation do not always exhibit smoothness [10, 11, 13], that is, their solutions may not be $C^{\infty}$. This raises the issue of what minimal regularity is necessary to ensure the well-definition of the associated one-forms.
Although the theory of PSS equations generally assumes smooth solutions, this requirement is often implicit. Exceptions are rare, with the work [37] standing out as a notable example. There, PSS surfaces derived from solutions to a specific Cauchy problem were considered, but the solutions in question were analytic ($C^{\omega}$), and hence smooth. Nevertheless, [37] also demonstrated the existence of a strip in which the one-forms are defined and the condition $\omega_{1}\wedge\omega_{2}\neq 0$ is satisfied.
It is worth recalling that, for a solution to define a first fundamental form of a surface, the associated one-forms must be linearly independent over simply connected domains in the plane [9, Theorem 4.39]. This requirement prompts further examination of how the geometry of the solution’s graph evolves, particularly with respect to the domain where a PSS structure can be meaningfully defined.
Solutions of the CH equation may develop singularities in finite time. For example, Constantin and Escher [11, Example 4.3] demonstrated that even smooth initial data may lead to wave breaking in finite time. The formation of such singularities is typically expressed through the blow-up of the slope of the solution.
The source of the problems above is precisely one: the space where a given solution is considered. From a geometric perspective, provided that the one-forms (2.0.2) are $C^{1}$ (see [23, Theorem 10-19, page 232] and also [23, Theorem 10-18, page 232]), we do have an abstract surface defined. Therefore, as long as we have a solution belonging to a function space respecting this constraint, the theory developed so far continues working.
With all of this in mind, we can look for minor modifications in the notion of PSS equations and generic solutions, e.g., see [20, 21].
From now on, we closely follows the ideas presented in [20]. Our presentation is more concerned with the geometric nature of the results and we avoid technicalities with functional analysis. However, we guide the reader for [20] to clarify certain aspects on the function spaces involved, such as the Sobolev spaces $H^{s}(\mathbb{R})$ and other related tools.
Definition 3.1.
[…]
  • the functions $f_{ij}$ are $C^{k}$ with respect to their arguments;
  • the forms (2.0.2) satisfy the structure equations (2.0.3) of a pseudospherical surface of Gaussian curvature ${\cal K}=-1$;
  • $\omega_{1}\wedge\omega_{2}\neq 0$.
If the function space is the space of $C^{\infty}$ functions, or it is not exactly or clearly defined, and no confusion is possible, we then say PSS equation instead of ${\cal B}-$PSS equation.
Remark 3.1.
The one-forms involved in Definition 2.1 are pull-backs of differential forms on jet spaces; see [30, Section II].
Definition 3.2.
(Generic solution) A function $u:U\rightarrow\mathbb{R}$ is called a generic solution for the ${\cal B}-$PSS equation (2.0.1) provided that:
  1. a)
    $u\in{\cal B}$;
  2. b)
    it is a solution of the equation;
  3. c)
    the one-forms (2.0.2) are $C^{k}$ on $U$;
  4. d)
    there exists at least a simply connected open set $\Omega\subseteq U$ such that $\omega_{1}\wedge\omega_{2}\big{|}_{p}\neq 0$, for each $p\in\Omega$.
Otherwise, $u$ is said to be non-generic.
Remark 3.2.
The condition $\omega_{1}\wedge\omega_{2}\big{|}_{p}\neq 0$ in Definition 3.2 means that
$\Big{(}f_{11}f_{22}-f_{12}f_{21}\Big{)}(p,u(p),u_{(1)}(p),\cdots,u_{(n)}(p))\neq 0.$
Additionally, all forms we will deal with herein have to be understood in a similar fashion to the above comment. For a better discussion, see [31, pages 77–78].
Lemma 3.1.
Assume that $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$ is a non-trivial solution of the CH equation (1.0.6). If $u_{x}\neq 0$ on some open set $\Omega$, then we cannot have $m=c\neq 0$, where $c$ is a constant.
Proof.
Suppose $m=c\neq 0$ is a constant on some open and non-empty set $\Omega$, then necessarily $m\neq 0$ and (1.0.6) implies (2.0.11). On the other hand, if $u$ is a solution of (1.0.6) such that $u_{x}=0$ on an open set $\Omega$, then $u_{x}=u_{xx}=u_{xxx}=u_{txx}=0$ and $u_{t}=0$ on $\Omega$, which, once substituted into (1.0.6), implies $u(x,t)=c$. ∎
Theorem 3.1.
Let $u_{0}\in H^{4}(\mathbb{R})$ be a non-trivial initial datum and $u$ be the corresponding solution of the CH equation subject to $u(x,0)=u_{0}(x)$. Then there exists a strip ${\cal S}=\mathbb{R}\times(0,T)$, uniquely defined by $u_{0}$, such that:
  1. a)
    the one-forms (2.0.8) are defined on ${\cal S}$;
  2. b)
    there exist at least two open and disjoint discs endowed with the structure of a $C^{1}$ PSS modelled by $C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$.
Proof.
Let $u_{0}\in H^{4}(\mathbb{R})$. From [10, Proposition 2.7] with $s=4$ we can find a number $T>0$, uniquely determined by $u_{0}$, and determine a unique function $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$ that is a solution of (1.0.5). The lifespan $T$ determined by the initial datum uniquely defines the strip mentioned in the statement of the theorem, whereas the solution $u$, granted by [10, Proposition 2.7], ensures that (2.0.9) are $C^{1}$ one-forms defined on ${\cal S}$.
Since $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$, we have $u_{x}\in H^{3}(\mathbb{R})$ for each fixed $t\in[0,T)$. Moreover, since the $H^{1}(\mathbb{R})$-norm of the solutions of the CH equation is invariant, and the fact that $u_{0}$ is non-trivial, imply that for any value of $t$ for which the solution exists, we have
$\|u(\cdot,t)\|_{H^{1}(\mathbb{R})}=\|u_{0}(\cdot)\|_{H^{1}(\mathbb{R})}>0,$ (3.0.1)
telling us that
$\lim_{|x|\rightarrow\infty}u(x,t)=0\quad\text{and}\quad\lim_{|x|\rightarrow\infty}u_{x}(x,t)=0.$ (3.0.2)
As a result, for each fixed $t\in[0,T)$, the function $x\mapsto u_{x}(x,t)$ has a minimum and a maximum. Let $a_{t}$ and $b_{t}$, with $a_{t}<b_{t}$, be points where these extrema are achieved (note that they are not necessarily unique). Let $I(t):=u_{x}(a_{t},t)$, $S(t):=u_{x}(b_{t},t)$ and $h(t):=I(t)S(t)$.
Claim. We claim that $h(t)<0$.
Let $a_{t}$ and $b_{t}$ be the distinct points we found earlier. Without loss of generality, we may assume $u_{x}(a_{t},t)<0<u_{x}(b_{t},t)$. By continuity, we can find $\epsilon>0$ such that $B_{\epsilon}(a_{t},t)$ and $B_{\epsilon}(b_{t},t)$ are disjoint, their union is contained in ${\cal S}$, and
$u_{x}\Big{|}_{B_{\epsilon}(b_{t},t)}>0\quad\text{and}\quad u_{x}\Big{|}_{B_{\epsilon}(a_{t},t)}<0.$
By Lemma 3.1, $m$ cannot be constant on $B_{\epsilon}(b_{t},t)$. As a result, for some $(y_{0},t_{0})\in B_{\epsilon}(b_{t},t)$ we have $m(y_{0},t_{0})\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$. Then, by continuity, we can find $\epsilon^{\prime}\in(0,\epsilon]$ such that $m(x,t)\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$, for all $(x,t)\in B_{\epsilon^{\prime}}(y_{0},t_{0})\subseteq B_{\epsilon}(b_{t},t)$. Let $B_{1}:=B_{\epsilon^{\prime}}(y_{0},t_{0})$. The same argument shows the existence of another disc $B_{2}\subseteq B_{\epsilon}(a_{t},t)$ where $m\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$.
Therefore, by the first part of the theorem, the one-forms (2.0.8) are defined on these discs, while the conditions above say that $\omega_{1}\wedge\omega_{2}\big{|}_{B_{1}}\neq 0$, $\omega_{1}\wedge\omega_{2}\big{|}_{B_{2}}\neq 0$, and they do not vanish at any point of $B_{1}\cup B_{2}$.
Let us now prove the claim by showing that we cannot have $h(t)\geq 0$. If we had $h(t_{0})\geq 0$ for some $t_{0}\in[0,T)$, then either $u_{x}(x,t_{0})\geq 0$ or $u_{x}(x,t_{0})\leq 0$ for all $x\in\mathbb{R}$, meaning that $u$ is either non-decreasing or non-increasing. Therefore, (3.0.2) implies $u(\cdot,t_{0})\equiv 0$, which contradicts (3.0.1). ∎
We have just shown that any non-trivial initial datum defines uniquely a strip ${\cal S}$ containing open sets that can be endowed with the structure of a PSS. Our task now is to look for singularities of the metric defined by the forms (2.0.8). By a singularity, we mean a point $p$ for which the metric is singular, in the sense that it is no longer a positive definite bilinear form. This situation occurs whenever the one-forms $\omega_{1}$ and $\omega_{2}$ are not linearly independent at $p$, that is, $\omega_{1}\wedge\omega_{2}\big{|}_{p}=0$.
Theorem 3.2.
Assume that $u_{0}\in H^{4}(\mathbb{R})$ is a non-trivial initial datum, $u$ the corresponding solution of (1.0.5), and let ${\cal S}$ be the strip determined in Theorem 3.1.
  1. a)
    If there exists some open set $\Omega\subseteq{\cal S}$ such that $\omega_{1}\wedge\omega_{2}\big{|}_{\Omega}=0$, then necessarily $u(x,t)=c$ for some $c\in\mathbb{R}$, $(x,t)\in\Omega$, and $\Omega$ is a proper subset of ${\cal S}$;
  2. b)
    For each $t\in(0,T)$, there exists at least one point $c_{t}\in\mathbb{R}$ such that $\omega_{1}\wedge\omega_{2}\big{|}_{(c_{t},t)}=0$.
Proof.
By Example 2.2, $u$ is a generic solution of the CH equation provided that $u_{x}=0$ or $m=const.$ For convenience, let us consider the CH equation in the form (1.0.6). If we assume that (2.0.10) holds on some open and non-empty set $\Omega$, then necessarily $m\neq 0$ and (1.0.6) implies (2.0.11). On the other hand, Lemma 3.1 implies that if $m$ is constant on some open set, then $u(x,t)$ coincides with $m$ within the same open set.
Let us show the existence of at least one point $c_{t}$ such that $u_{x}(c_{t},t)=0$ for each $t\in[0,T)$. Let $h$, $a_{t}$ and $b_{t}$ be as in the proof of Theorem 3.1. The two points $a_{t}$ and $b_{t}$ cannot be the same and the intermediate value theorem ensures the existence of $c_{t}\in(a_{t},b_{t})$ such that $u_{x}(c_{t},t)=0$.
Finally, let us prove that $\Omega$ must be a proper subset of ${\cal S}$. In fact, if we had $\Omega={\cal S}$, then we would be forced to conclude that $c=0$ from (3.0.2), which would then lead to a contradiction with (3.0.1). ∎
4 Final remarks
Remark 4.1.
The investigation of PSS equations essentialy begun with the papers by Sasaki [38] and Chern and Tenenblat [8]. Both papers were motivated by the geometric nature of certain integrable equations in the sense describled in [1].
Remark 4.2.
The paper by Chern and Tenenblat [8], although motivated by Sasaki [38] started systematic geometric studies of integrable systems from the point of view of two-dimensional Riemannian manifolds. In consequence, nowadays we know that PSS equations are not only restricted to AKNS equations. In fact, equations with no integrable properties (in any sence) can be PSS equations, e.g., see [5] for some examples of non-integrable equations that are PSS equations.
Remark 4.3.
E. Reyes has made fundamental contributions in the study of PSS equations and integrable systems [3, 30, 31, 32, 33, 34, 35]. In particular, he proved that the CH equation is geometrically integrable [32]. In [33] the connections between PSS equations and integrability are explored, while in [35] a nice review on PSS in light of the legacy by Chern and Tenenblat contributions is presented.
Remark 4.4.
A topic not treated in this paper is the question of immersion, into the Euclidean space, of the abstract surfaces determined by the solutions of PSS equations. It is well known that PSS can be (locally) immersed into the three-dimensional space. In particular, there are works studying the immersion of surfaces determined by equations, see [6, 21, 24, 25, 26, 30] and references therein.
Remark 4.5.
The results reported in [6] say that the immersion of a PSS determined by a generic solution of the CH equation must depend on an infinite number of derivatives of the solution. However, the results established in [20], see also [16, 21, 22] tell us that the manifold may have a metric with finite regularity (see Theorem 3.1). This is, however, not a contradiction. In [6] it is assumed that the solutions are $C^{\infty}$ and then, the natural locus for all the involved functions and forms are the jet space, eventually an infinite jet space. The results in [20] are obtained in a context of finite jets, meaning that a second fundamental form cannot have arbitrary derivatives. In particular, the results in [20] open doors for us to investigate second fundamental forms having a different dependence on the field variables.
Acknowledgements
I would like to thank Professor Tiago Macedo for his condition for me to contribute to CM with this review paper. I am also thankful to CNPq (grant nº 310074/2021-5) and FAPESP (grant nº 2024/01437-8) for financial support.